HETAV PATEL

Grade 11,

a body covers a half distance v1 and remain half with v2 between two station find average velocitt

a body covers a half distance v1 and remain half with v2 between two station find average velocitt

Grade:11

1 Answers

Anurag
6 years ago
Let the total distance be 2d.Velocity for half distance =v1Time taken to cover d distance=t1=d/v1 Velocity for other half=d/v2Time taken to cover this distance=t2=d/v2Total time=t1+t2=d/v1+d/v2=d (1/v1+1/v2)=d{ (v2+v1)/v1v2}Now avg velocity=Displacement/Time=2d/d{ (v2+v1)/v1v2}=2v1v2/v2+v1.